OpenStudy (anonymous):

Two hundred tickets for the school play were sold. Tickets cost $2 for students and $3 for adults. The total amount collected was $490. How many student tickets were sold?

hero (hero):

Two hundred tickets for the school play were sold. s + a = 200 Tickets cost $2 for students and $3 for adults. The total amount collected was $490. 2s + 3a = 490 How many student tickets were sold? Solve for s

OpenStudy (anonymous):

Please Don't give me any equations. That's too hard.

OpenStudy (anonymous):

REALLY. with the equations.

hero (hero):

I can post the solution if you want

OpenStudy (anonymous):

And THIS is why you're blocked. Dude loosen up!! Through the old rule book out the window, no one pays attention to it anyway.

OpenStudy (anonymous):

Answer please?

hero (hero):

I can still report you if you want that

hero (hero):

I already showed you the Code of Conduct

OpenStudy (anonymous):

Oh.. My.. Freaking.. Gosh bro.

OpenStudy (anonymous):

Like i said, no one reads that.

hero (hero):

If you say so

OpenStudy (anonymous):

& what are you gonna report me for exactly "mr. i'm a moderator person" IN CASE YOU HAVEN'T NOTICED, i haven't done anything wrong.

hero (hero):

Neither have I, but you blocked me anyway

hero (hero):

s + a = 200 2s + 3a = 490 a = 200 - s a = (490 - 2s)/3 a = a 3(200 - s) = 490 - 2s 600 - 3s = 490 - 2s 600 - 490 = 3s - 2s 110 = s

OpenStudy (anonymous):

Fine you want me to unfreaking block you? Give me 1 GOOD reason, and i will.

hero (hero):

I just posted the solution to your question

OpenStudy (anonymous):

THAT is not a solution. That's a big long thing that i'm having to work out.

hero (hero):

110 student tickets were sold

OpenStudy (anonymous):

Mkay.

hero (hero):

@KNicksfan4life , I don't appreciate your dry humor.

OpenStudy (anonymous):

He didn't even post anything.

hero (hero):

He posted something then quickly deleted it.

hero (hero):

Have a nice day Jesse and Knicks Fan. Go Jeremy Lin? lol